If P and V are the power and potential of device,

the power consumed with a supply potential V

is


1
Expert's answer
2021-06-29T05:44:01-0400

Let R be the Resistance of the device

"R=\frac{V^2}{P}\\"

The power consumed with a supply of V is

"P_c=\frac{V^2}{R}\\"

Substitute R in the equation above

"P_c=\frac{V^2}{\frac{V^2}{P}}\\
P_c=P"